OpenStudy (anonymous):

Solve. 81^x = 27^x + 2 My answer: x = 2 x = 1 x = 2 x = 5 x = 6

OpenStudy (anonymous):

ok that was wrong lets try again \[3^{4x}=3^{3x+6}\]

OpenStudy (anonymous):

how would I solve that?

OpenStudy (anonymous):

since \(3^4=81\) and \(27=3^3\)

OpenStudy (anonymous):

since we arranged it so the bases are the same, then the exponents must be the same that was the point of making the bases equal solve \[4x=3x+6\]

OpenStudy (anonymous):

oh, okay, I see now!!!

OpenStudy (anonymous):

ok good!

OpenStudy (anonymous):

I got 6 as my answer on my paper

OpenStudy (anonymous):

yes that should be right you can check it if you have a calculator that is check that \(81^6=27^8\) it is too much to do by hand, they are huge numbers

OpenStudy (anonymous):

okay i will right now

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw